Simple Plan Enterprisee uses a perlodic Inventory system. Its records showed the foliowng: Iriventory, Decenber 31, using FIFPlease answer with detail
Answer #1
FIFO
Units Rate p.u ($) Cost of Goods Available for Sale Sold out of Left in Ending Inventory Cost of Inventory($) Cost Of Goods Sold($)
a b c d e=a+d f=d*e g=-d*b
Beginning Inventory 44 $             19 $                 836 -44 0 $                       -   $                        836
Jan, 9 56 $             20 $              1,120 -56 0 $                       -   $                     1,120
Jan, 20 106 $             21 $              2,226 -48 58 $                  1,218 $                     1,008
Closing Inventory 58
Units Sold 148 $              4,182 -148 58 $                  1,218 $                     2,964
LIFO
Units Rate p.u ($) Cost of Goods Available for Sale Sold out of Left in Ending Inventory Cost of Inventory($) Cost Of Goods Sold($)
a b c d e=a+d f=d*e g=-d*b
Beginning Inventory 44 $             15 $                 660 -30 14 $                     210 $                        450
Jan, 9 56 $             20 $              1,120 -56 0 $                       -   $                     1,120
Jan, 20 106 $             21 $              2,226 -62 44 $                     924 $                     1,302
Closing Inventory 58
Units Sold 148 $              4,006 -148 58 $                  1,134 $                     2,872
1
FIFO LIFO
Number of Goods Available for Sale 206 206
Cost of Goods Available for Sale $ 4,182 $        4,006
Cost of Ending Inventory $ 1,218 $        1,134
Cost of Goods Sold $ 2,964 $        2,872
2
FIFO LIFO
Cost of Goods sold a $        2,964 $              2,872
Beginning Inventory b $           836 $                 660
Closing Inventory c $        1,218 $              1,134
Average Inventory d=b+c/2 $        1,027 $                 897
Inventory Turnover Ratio e=a/d 2.89 3.20